If the range of the function f(x)= 5x − 3 is {-13, -8, -3, 12, 17}, what is its domain?
Answer by Fombitz(32388)

Work backwards from the range.
5x-3=-13
5x=-10
x=-2
.
.
5x-3=-8
5x=-5
x=-1
.
.
5x-3=-3
5x=0
x=0
.
.
5x-3=12
5x=15
x=3
.
.
5x-3=17
5x=20
x=4
.
.
Domain: {-2,-1,0,3,4}